Replace the loading system by an equivalent resultant force and couple moment at 0.X 2016 aft В F = (21-4j +56)1b -24 || 1-244t-aft Mo 10 16.67 Us So16 Replace the loading system by on equivalent resultant fo
Answer #1

N to-ordinates of various points: A(-2,6, 0), 4+ B(-4, 6, o), ft c(-1,0, 2), At D(0, 6, o), ft Forces F = (22-49 +5%), u a =moment due to WA WA - j k о е = - 300 ? - 100 lbft 0 0 - 50 moment due to Fn = 7x Fr MB - MB = j k -6 = 1(-240) - § (-160) +